# Don’t Break the Cache: An Evaluation of Prompt Caching for Long-Horizon Agentic Tasks

### v1

@@ Abstract
Recent advancements in Large Language Model (LLM) agents have enabled complex multi-turn agentic tasks requiring extensive tool calling, where conversations can span dozens of API calls with increasingly large context windows. However, although major LLM providers offer prompt caching to reduce cost and latency, its benefits for agentic workloads remain underexplored in the research literature. To our knowledge, no prior work quantifies these cost savings or compares caching strategies for multi-turn agentic tasks. We present a comprehensive evaluation of prompt caching across three major LLM providers (OpenAI, Anthropic, and Google) and compare three caching strategies, including full context caching, system prompt only caching, and caching that excludes dynamic tool results. We evaluate on DeepResearchBench, a multi-turn agentic benchmark where agents autonomously execute real-world web search tool calls to answer complex research questions, measuring both API cost and time to first token (TTFT) across over 500 agent sessions with 10,000-token system prompts. Our results demonstrate that prompt caching reduces API costs by 45-80% and improves time to first token by 13-31% across providers. We find that strategic prompt cache block control, such as placing dynamic content at the end of the system prompt, avoiding dynamic traditional function calling, and excluding dynamic tool results, provides more consistent benefits than naive full-context caching, which can paradoxically increase latency. Our analysis reveals nuanced variations in caching behavior across providers, and we provide practical guidance for implementing prompt caching in production agentic systems.
@@ 1 Introduction
Recent advancements in Large Language Model (LLM) agents have enabled complex, long-horizon agentic tasks that require extensive tool calling across multi-turn conversations (Ji, 2025). Through function calling, LLM agents can invoke external APIs, execute web searches, interact with databases, and perform domain-specific actions on behalf of users. As these agentic workloads grow in complexity, conversations can span dozens of API calls with context windows accumulating tens of thousands of tokens, leading to significant costs and latency overhead. To address this, major LLM providers including OpenAI, Anthropic, and Google now offer prompt caching, a feature that reuses previously computed key-value (KV) tensors from attention layers to avoid redundant computation on repeated prompt prefixes (OpenAI, 2026; Anthropic, 2026; Google Cloud, 2026a).
Figure 1: Prompt caching benefits (best cache mode per model). Percentage reduction in API cost and time to first token (TTFT) relative to a no-cache baseline. For each model, results correspond to the best-performing cache strategy. Asterisks denote statistically significant TTFT improvements ($p<0.05$).
@@
In this paper, we present the first comprehensive evaluation of prompt caching strategies for long-horizon agentic tasks across three major LLM providers (OpenAI, Anthropic, and Google) using four flagship models (Figure 1). We compare three caching strategies, including full context caching, system prompt only caching, and caching that excludes dynamic tool results. We evaluate on DeepResearchBench (Du et al., 2025), a multi-turn agentic benchmark where agents autonomously execute web search tool calls to answer complex research questions. Our evaluation spans over 500 agent sessions with 10,000-token system prompts, measuring both API cost and time to first token (TTFT) across all conditions.
@@
Prompt caching delivers substantial and consistent cost savings across all providers: All four models tested show statistically significant cost reductions when prompt caching is enabled. Cost savings range from 45% to 80% across providers. These savings are consistent across all three caching strategies, demonstrating that prompt caching provides reliable cost benefits regardless of the specific caching approach employed.
@@ 2.1 KV Cache and LLM Inference
As context windows have grown from thousands to millions of tokens, KV cache management has become a critical bottleneck in LLM serving (Shi et al., 2024). The memory footprint of KV caches scales linearly with sequence length and batch size, often consuming more GPU memory than the model weights themselves for long-context workloads. This has motivated extensive research on KV cache optimization, including memory management techniques such as PagedAttention (Kwon et al., 2023), which applies paging-style memory management to reduce fragmentation and waste, achieving 2-4x throughput improvements. Other approaches focus on KV cache compression through selective retention of important tokens (Ge et al., 2023), storage-compute tradeoffs that balance recomputation against cache loading (Jin et al., 2024), and shared prefix optimization for high-throughput inference (Juravsky et al., 2024; Wu and others, 2024; Sun et al., 2025; Zhou et al., 2024).
@@ 2.2 Prompt Caching in Provider APIs
Major LLM providers have implemented prompt caching with varying approaches. OpenAI offers automatic prompt caching on GPT-4o and newer models, where caching activates automatically for prompts exceeding a minimum token threshold, with cache hits occurring only for exact prefix matches (OpenAI, 2026, 2024a). Anthropic provides developer-controlled caching through explicit cache breakpoints, allowing users to specify which portions of their prompt should be cached, with configurable time-to-live (TTL) options (Anthropic, 2026, 2025c). Google offers both implicit caching, which activates automatically with no guaranteed cost savings, and explicit context caching, where developers create and reference caches with guaranteed discounts (Google Cloud, 2026a; Kilpatrick, 2025).
@@ 2.3 Agentic Workloads and Context Engineering
Recent advances in LLM agents have enabled complex, long-horizon tasks that extend far beyond single-turn question answering. Modern agentic applications including deep research assistants, coding agents such as Claude Code and Cursor, and autonomous task completion systems like Manus routinely execute 30-50 or more tool calls within a single session (Ji, 2025; Du et al., 2025; Mialon et al., 2023; Zhou et al., 2023; Drouin et al., 2024; Wei et al., 2025). Each tool call adds content to the conversation context, including the tool invocation, execution results, and the model’s subsequent reasoning, causing context windows to grow rapidly throughout the session.
This growth presents unique challenges for prompt caching (Guan et al., 2026; Laban et al., 2025). Recent work has proposed solutions for multi-turn caching scenarios (Jeong and Ahn, 2025; Yan et al., 2025). Unlike static question-answering scenarios where prompts are largely predetermined, agentic workloads feature dynamic, session-specific content that accumulates unpredictably. Tool results often contain user-specific data that will not benefit other sessions, and the interleaving of static system prompts with dynamic tool outputs complicates cache reuse. Context engineering strategies have emerged to manage these challenges, including treating external storage as extended memory and carefully structuring prompts to maximize cache efficiency (Ji, 2025; Lumer et al., 2025a). However, the effectiveness of prompt caching across different caching strategies in agentic workloads has not been comprehensively evaluated. Our work addresses this gap by measuring cost and latency benefits across controlled caching strategies on a multi-turn agentic benchmark.
@@ 3.1 Experimental Setup
We evaluate prompt caching across three major LLM providers: OpenAI, Anthropic, and Google. For each provider, we select a flagship model: GPT-4o and GPT-5.2 from OpenAI, Claude Sonnet 4.5 from Anthropic, and Gemini 2.5 Pro from Google. These models represent the current state-of-the-art for agentic workloads and all support prompt caching through their respective APIs.
We use DeepResearchBench (Du et al., 2025) as our evaluation benchmark, a multi-turn agentic benchmark where agents autonomously execute web search tool calls to answer complex research questions. We selected this benchmark over alternatives such as other deep research benchmarks (FutureSearch: Bosse et al., 2025; Li et al., 2025) due to its focus on tool-intensive agentic workflows and real-world 100 PhD-level research tasks, each meticulously crafted by domain experts across 22 distinct fields. We implement our research agent using Deep Agents (LangChain, 2025), one of various open source libraries for creating long-running agents (Anthropic, 2025a; OpenAI, 2025; Google, 2025; Microsoft, 2023a, 2026, b; CrewAI, 2023; LlamaIndex, 2022; Hugging Face, 2024; OpenAI, 2024b; Agno, 2026). Each agent session begins with a research question and the agent iteratively calls a web search tool to gather information before synthesizing a comprehensive response. This benchmark reflects realistic agentic workloads where context windows grow dynamically through tool invocations and results.
@@ 4.1 Overall Results
Table 1: Prompt caching benefits by model using the best-performing cache mode for each. Cost savings and TTFT improvement are relative to the no-cache baseline. Bold indicates highest value per metric.
@@ 4.2 Cost Reduction
Figure 3: Normalized cost and time to first token (TTFT) distributions by model and cache strategy. The no-cache baseline is normalized to 100% and lower values indicate better performance. Cost reductions are consistent across cache strategies, while TTFT improvements vary significantly, with full-context caching sometimes underperforming more selective strategies.
@@ 5.1 Strategic Cache Boundary Control
The most effective strategy is to ensure that only stable, reusable content is cached. In agentic applications, the system prompt is the most stable component, containing agent instructions, tool definitions, and persona guidelines that remain constant across sessions. Conversation history, tool calls, and tool results are inherently dynamic and session-specific, making them poor candidates for cross-session caching. Practitioners should avoid including dynamic values in the system prompt itself. Common patterns that inadvertently break the cache include timestamps, datetime strings, session identifiers, or user-specific information embedded in the system prompt. If such dynamic information is necessary, it should be placed at the end of the system prompt to maximize the cacheable prefix. This ensures that the majority of the system prompt benefits from cache hits while only the dynamic suffix requires recomputation (Ji, 2025).
Similarly, dynamic function calling can break the cache when tool definitions change between requests. Modern agentic systems increasingly leverage dynamic tool discovery and registration through protocols such as the Model Context Protocol (MCP) (Model Context Protocol, 2026), where available tools may vary based on connected servers or runtime context (Lumer et al., 2025a). When tool definitions are included in the prompt, any change to the available tool set invalidates the cached prefix. A practical strategy is to maintain a fixed set of general-purpose, reusable functions (such as code execution, file operations, and bash/shell commands), while implementing dynamic capabilities through code generation rather than traditional function calling (Ji, 2025; Wang et al., 2024; Jones and Kelly, 2025; Anthropic, 2025b; Varda and Pai, 2025; Hacker News, 2025). Those prior methods of dynamic function calling, while achieving strong retrieval and execution accuracy, can prevent prompt caching usage (Lumer et al., 2024, 2025b; Chen et al., 2024; Zheng et al., 2024; Chen et al., 2024; Wu et al., 2024).
@@ 5.2 Tool Call Caching Considerations
For long-running agentic sessions with 30-50 or more tool calls, practitioners may consider caching tool calls and results to further reduce costs. However, this approach involves tradeoffs. Cache creation incurs a cost and latency overhead on the first request, which is only amortized if subsequent requests benefit from cache reads. For tool calls that produce highly variable results or that are unlikely to be repeated, caching provides no benefit and may introduce unnecessary overhead.
Furthermore, common context management strategies in agentic systems can interact poorly with tool call caching. Techniques such as summarizing or pruning old tool calls to manage context length (Ji, 2025) inherently modify the conversation history, breaking any cached representations of that content. If an application employs such strategies, caching tool calls becomes counterproductive. The emerging pattern for agentic applications is to maintain a large, stable system prompt that benefits from caching, while treating tool calls and results as dynamic content that may be summarized, pruned, or otherwise managed throughout the session.
@@ 5.3 Provider Implementation Variability
Our results also reflect natural variance in API response times due to factors including server load, geographic distribution, and infrastructure differences across providers. When evaluating prompt caching benefits, practitioners should conduct experiments representative of their specific workloads and usage patterns rather than relying solely on published benchmarks. Practitioners should also be aware of security considerations, as recent work has demonstrated that prompt caching can introduce timing side-channels that may leak information about cached content (Wu et al., 2025; Gu et al., 2025).
@@ 6 Conclusion
Recent advancements in Large Language Model (LLM) agents have enabled complex multi-turn agentic tasks requiring extensive tool calling, where conversations can span dozens of API calls with increasingly large context windows. While major LLM providers offer prompt caching to reduce costs and latency, the benefits of these features remain under-explored in the research literature. In this work, we present a comprehensive evaluation of prompt caching across three major LLM providers (OpenAI, Anthropic, and Google) using four flagship models. We compare three caching strategies, including full context caching, system prompt only caching, and caching that excludes dynamic tool results. We evaluate on DeepResearchBench, a multi-turn agentic benchmark where agents autonomously execute web search tool calls to answer complex research questions, measuring both API cost and time to first token (TTFT) across over 500 agent sessions with 10,000-token system prompts. Our results demonstrate that prompt caching reduces API costs by 45-80% and improves time to first token by 13-31% across providers. We find that strategic cache boundary control, such as excluding dynamic tool results, provides more consistent benefits than naive full-context caching, which can paradoxically increase latency. As agentic systems continue to trend toward longer-running sessions with dozens of tool calls, strategic prompt caching becomes critical for reducing operational costs and improving user experience, and our findings provide practitioners with guidance for implementing prompt caching in production agentic systems.

### v2

@@ Abstract
Recent advancements in Large Language Model (LLM) agents have enabled complex multi-turn agentic tasks requiring extensive tool calling, where conversations can span dozens of API calls with increasingly large context windows. However, although major LLM providers offer prompt caching to reduce cost and latency, its benefits for agentic workloads remain underexplored in the research literature. To our knowledge, no prior work quantifies these cost savings or compares caching strategies for multi-turn agentic tasks. We present a comprehensive evaluation of prompt caching across three major LLM providers (OpenAI, Anthropic, and Google) and compare three caching strategies, including full context caching, system prompt only caching, and caching that excludes dynamic tool results. We evaluate on DeepResearch Bench, a multi-turn agentic benchmark where agents autonomously execute real-world web search tool calls to answer complex research questions, measuring both API cost and time to first token (TTFT) across over 500 agent sessions with 10,000-token system prompts. Our results demonstrate that prompt caching reduces API costs by 41-80% and improves time to first token by 13-31% across providers. We find that strategic prompt cache block control, such as placing dynamic content at the end of the system prompt, avoiding dynamic traditional function calling, and excluding dynamic tool results, provides more consistent benefits than naive full-context caching, which can paradoxically increase latency. An ablation study across prompt sizes (500-50,000 tokens) and tool call counts (3-50) demonstrates universal linear cost and TTFT benefits, after the provider caching token minimum, and reveal provider-specific strategy discrepancies across variants. We provide nuanced discussion and guidance for implementing prompt caching in production agentic systems.
@@ 1 Introduction
Recent advancements in Large Language Model (LLM) agents have enabled complex, long-horizon agentic tasks that require extensive tool calling across multi-turn conversations (Ji, 2025). Through function calling, LLM agents can invoke APIs, execute web searches, interact with databases, and perform domain-specific actions on behalf of users. As these agentic workloads grow in complexity, conversations can span dozens of API calls with context windows accumulating tens of thousands of tokens, leading to significant costs and latency overhead. To address this, major LLM providers including OpenAI, Anthropic, and Google offer prompt caching, a feature that reuses previously computed key-value (KV) tensors from attention layers to avoid redundant computation on repeated prompt prefixes (OpenAI, 2026; Anthropic, 2026; Google Cloud, 2026a).
Figure 1: Prompt caching benefits (best cache mode per model). Percentage reduction in API cost and time to first token (TTFT) relative to a no-cache baseline. Asterisks denote statistically significant TTFT improvements ($p<0.05$).
@@
In this paper, we present the first comprehensive evaluation of prompt caching strategies for long-horizon agentic tasks across three LLM providers (OpenAI, Anthropic, and Google) using four flagship models (Figure 1). We compare three caching strategies, including full context caching, system prompt only caching, and caching that excludes dynamic tool results. We evaluate on DeepResearch Bench (Du et al., 2025), a multi-turn agentic benchmark where agents autonomously execute web search tool calls to answer research questions. Our evaluation spans 500 agent sessions with 10,000-token system prompts, measuring both API cost and time to first token (TTFT) across all conditions.
@@
Prompt caching delivers substantial and consistent cost savings across all providers: All four models tested show statistically significant cost reductions when prompt caching is enabled. Cost savings range from 41% to 80% across providers. These savings are consistent across all three caching strategies, demonstrating that prompt caching provides reliable cost benefits regardless of the specific caching approach employed.
@@ 2.1 KV Cache and LLM Inference
This has motivated extensive research on KV cache optimization, including memory management techniques such as PagedAttention (Kwon et al., 2023), which applies paging-style memory management to reduce fragmentation and waste. Other approaches focus on KV cache compression through selective retention of important tokens (Ge et al., 2023), storage-compute tradeoffs (Jin et al., 2024), and shared prefix optimization for high-throughput inference (Juravsky et al., 2024; Wu and others, 2024; Sun et al., 2025; Zhou et al., 2024).
@@ 2.2 Prompt Caching in Provider APIs
Major LLM providers have implemented prompt caching with varying approaches. OpenAI offers automatic prompt caching on GPT-4o and newer models, where caching activates automatically for prompts exceeding a minimum token threshold, with cache hits occurring only for exact prefix matches (OpenAI, 2026, 2024a). Anthropic provides developer-controlled caching through explicit cache breakpoints, allowing users to specify which portions of their prompt should be cached, with configurable time-to-live (TTL) options (Anthropic, 2026, 2025b). Google offers both implicit caching, which activates automatically with no guaranteed cost savings, and explicit context caching, where developers create and reference caches with guaranteed discounts (Google Cloud, 2026a; Kilpatrick, 2025).
@@ 2.3 Agentic Workloads and Context Engineering
Recent advances in LLM agents have enabled complex, long-horizon tasks that extend far beyond single-turn question answering. Modern agentic applications including deep research assistants, coding agents such as Claude Code and Cursor, and autonomous task completion systems like Manus routinely execute 30-50 or more tool calls within a single session (Ji, 2025; Du et al., 2025; Mialon et al., 2023; Zhou et al., 2023; Drouin et al., 2024; Wei et al., 2025). In such workflows, each tool call adds content to the conversation context, including the tool invocation, execution results, and the model’s subsequent reasoning, causing context windows to grow rapidly throughout the session.
This growth presents challenges for prompt caching (Guan et al., 2026; Laban et al., 2025). Recent work has proposed solutions for multi-turn caching scenarios (Jeong and Ahn, 2025; Yan et al., 2025). Unlike static question-answering scenarios where prompts are largely predetermined, agentic workloads feature dynamic, session-specific content that accumulates unpredictably. Tool results often contain user-specific data that will not benefit other sessions, and the interleaving of static system prompts with dynamic tool outputs complicates cache reuse. Context engineering strategies have emerged to manage these challenges, including treating external storage as extended memory and structuring prompts to maximize cache efficiency (Ji, 2025; Lumer et al., 2025a). However, the effectiveness of prompt caching across different caching strategies in agentic workloads has not been comprehensively evaluated. Our work addresses this gap by measuring cost and latency benefits across controlled caching strategies on a multi-turn agentic benchmark.
@@ 3.1 Experimental Setup
We evaluate prompt caching across three major LLM providers: OpenAI, Anthropic, and Google. For each provider, we select flagship models: GPT-5.2 from OpenAI, Claude Sonnet 4.5 from Anthropic, and Gemini 2.5 Pro from Google. We additionally include GPT-4o to examine caching behavior across model generations. All selected models support prompt caching through their respective APIs.
We use DeepResearch Bench (Du et al., 2025) as our evaluation benchmark, a multi-turn agentic benchmark where agents autonomously execute web search tool calls to answer complex research questions. We selected this benchmark over alternatives such as other deep research benchmarks (FutureSearch: Bosse et al., 2025; Li et al., 2025) due to its focus on tool-intensive agentic workflows and real-world 100 PhD-level research tasks, each crafted by domain experts across 22 fields. We implement our research agent using Deep Agents (LangChain, 2025), one of various open source libraries for creating long-running agents (Anthropic, 2025a; OpenAI, 2025; Google, 2025; Microsoft, 2023a, 2026, b; CrewAI, 2023; LlamaIndex, 2022; Hugging Face, 2024; OpenAI, 2024b; Agno, 2026). Each agent session begins with a research question and the agent iteratively calls a web search tool to gather information before synthesizing a comprehensive response. This benchmark reflects realistic agentic workloads where context windows grow dynamically through tool invocations and results.
@@ 4.1 Overall Results
Table 1: Prompt caching benefits using the best-performing cache mode for each. Cost savings and TTFT improvement are relative to the no-cache baseline. Bold indicates highest value per metric.
@@ 4.2 Cost Reduction
Figure 3: Normalized cost and TTFT distributions by model and cache strategy. Baseline = 100%, lower is better. Cost savings are consistent across strategies - TTFT varies, with full-context caching sometimes underperforming selective strategies.
@@ 5.1 Strategic Cache Boundary Control
The most effective strategy is to ensure that only stable, reusable content is cached. In agentic applications, the system prompt is the most stable component, containing agent instructions, tool definitions, and persona guidelines that remain constant across sessions. Conversation history, tool calls, and tool results are dynamic and session-specific, making them poor candidates for cross-session caching. Practitioners should avoid including dynamic values in the system prompt itself. Common patterns that inadvertently break the cache include timestamps, datetime strings, session identifiers, or user-specific information embedded in the system prompt. If such dynamic information is necessary, it should be placed at the end of the system prompt to maximize the cacheable prefix. This ensures that the majority of the system prompt benefits from cache hits while only the dynamic suffix requires recomputation (Ji, 2025).
Similarly, dynamic function calling can break the cache when tool definitions change between requests. Modern agentic systems increasingly leverage dynamic tool discovery through protocols such as the Model Context Protocol (MCP) (Model Context Protocol, 2026), where available tools may vary based on connected servers or runtime context (Lumer et al., 2025a). When tool definitions are included in the prompt, any change to the available tool set invalidates the cached prefix. A practical strategy is to maintain a fixed set of general-purpose, reusable functions while implementing dynamic capabilities through code generation rather than traditional function calling (Ji, 2025; Wang et al., 2024). Prior methods of dynamic function calling can prevent prompt caching usage (Lumer et al., 2024, 2025b).
@@ 5.2 Tool Call Caching Considerations
For long-running agentic sessions with 30-50+ tool calls, practitioners may consider caching tool calls and results to reduce costs. However, cache creation incurs overhead on the first request, which is only amortized if subsequent requests benefit from cache reads. For tool calls that produce variable results or are unlikely to be repeated, caching provides no benefit and may introduce overhead.
Common context management strategies can interact poorly with tool call caching. Techniques such as summarizing or pruning old tool calls (Ji, 2025) break cached representations, making tool call caching counterproductive. The emerging pattern is to maintain a stable system prompt that benefits from caching while treating tool calls as dynamic content that may be managed throughout the session.
@@ 5.3 Provider Implementation Variability
Our results reflect natural variance in API response times due to factors including server load, geographic distribution, and infrastructure differences across providers. When evaluating prompt caching benefits, practitioners should conduct experiments representative of their workloads and usage patterns rather than relying solely on published benchmarks. Practitioners should be aware of security considerations, as recent work has demonstrated that prompt caching can introduce timing side-channels that may leak information about cached content (Wu et al., 2025; Gu et al., 2025).
6 Ablation Study
To understand how prompt caching benefits scale with workload characteristics, we conduct an ablation study examining two key dimensions: prompt size and tool call count. These experiments isolate the factors that drive caching effectiveness and inform practitioners which workload characteristics benefit most from prompt caching. Figure 4 presents ablation results across three models and cache strategies.
Figure 4: Ablation results by prompt size (left) and tool count (right). Rows: GPT-4o, GPT-5.2, Claude Sonnet 4.5. Cost savings (columns 1, 3) are consistent across strategies vs. baseline (red). TTFT (columns 2, 4) varies more, below caching thresholds and at high counts.
6.1 Ablation by Prompt Size
We evaluate prompt caching across six prompt sizes: 500, 2,000, 5,000, 10,000, 20,000, and 50,000 tokens. For each size and model, we measure median cost and time to first token (TTFT) across all cache strategies. Notably, the 500-token condition falls below the minimum threshold required for prompt caching to activate (1,024 tokens for OpenAI and Anthropic, 4,096 tokens for Google; see Table 4), serving as a control condition where caching cannot provide benefits.
As shown in the left columns of Figure 4, cost savings scale linearly with prompt size across all models and are universally positive at every prompt size tested. At 50,000 tokens, GPT-5.2 achieves 89% cost savings (from \$0.253 to \$0.029), Claude Sonnet 4.5 achieves 88% savings (from \$0.667 to \$0.080), and GPT-4o achieves 54% savings (from \$0.414 to \$0.192). TTFT improvements are also most pronounced at large context sizes, with GPT-4o showing 60% improvement (from 4,290ms to 1,699ms) at 50,000 tokens.
At smaller prompt sizes (500-2,000 tokens), cost savings remain positive but modest, typically ranging from 10-45%. However, TTFT results at small prompt sizes reveal an important nuance: at 500 tokens, which falls below the minimum caching threshold, GPT-4o, GPT-5.2, and Claude Sonnet 4.5 show TTFT regressions of 10-18%. This is expected behavior, as the caching mechanism cannot activate below the threshold, and any observed latency differences reflect server variance rather than caching benefits. At prompt sizes above the threshold (2,000+ tokens), TTFT improvements become positive and scale with prompt size.
6.2 Ablation by Tool Count
We evaluate prompt caching across five tool call counts: 3, 5, 10, 20, and 50 tool calls per session. For each configuration, we measure median cost and TTFT across cache strategies. The results (Figure 4) demonstrate that cost savings remain consistent regardless of the number of tool calls in a session. GPT-5.2 maintains approximately 77-81% cost savings across all tool counts, while GPT-4o achieves 42-53% savings. This consistency indicates that the number of tool calls does not substantially affect the cost benefits of prompt caching. TTFT results for the tool count ablation show greater variance compared to the main study. While GPT-4o shows consistent TTFT improvements across all tool counts (16-36%), other models exhibit mixed results. Claude Sonnet 4.5 shows diminishing TTFT returns at higher tool counts (from 19% improvement at 3 tools to 5% at 50 tools). This increased variance likely reflects server load fluctuations during the ablation experiments, as the main study (Section 4) showed more consistent TTFT improvements with lower variance.
6.3 Discussion
Cost savings are universally positive. Across all models, prompt sizes, and tool counts tested, prompt caching consistently reduces API costs. This finding provides practitioners with confidence that enabling prompt caching will reliably reduce costs regardless of workload characteristics.
Prompt size drives caching benefits more than tool count. Cost savings scale linearly with prompt size (from 10-45% at 500 tokens to 54-89% at 50,000 tokens), while remaining stable across tool counts (typically within 10 percentage points). This indicates the cacheable prefix length, primarily determined by the system prompt, is the dominant factor in caching effectiveness. Practitioners should focus on maximizing system prompt size within architectural constraints rather than optimizing around tool call patterns.
TTFT improvements require meeting minimum thresholds. The TTFT regressions at 500 tokens highlight the importance of meeting provider-specific thresholds (Table 4). When prompts fall below these thresholds, caching cannot activate. The ablation study exhibits higher TTFT variance than the main study, reflecting the challenges of measuring latency in production API environments where server load and infrastructure changes introduce noise. Practitioners should conduct representative experiments for their workloads rather than relying on published benchmarks.
@@ 7 Conclusion
We present the first comprehensive evaluation of prompt caching for long-horizon agentic tasks across three major LLM providers. Our results demonstrate that prompt caching reduces API costs by 41 to 80% and improves TTFT by 13 to 31%. Strategic cache boundary control, such as caching only system prompts while excluding dynamic tool results, provides more consistent benefits than naive full context caching, which can paradoxically increase latency. These findings provide actionable guidance for deploying prompt caching in production agentic systems.